Find: (7/3) ÷ 2 Find: (7/3) ÷ 2 Find: (7/3) ÷ 2 Find: (7/3) ÷ 2

Find:
 (7/3) ÷ 2
Find:
 (7/3) ÷ 2
Find:
 (7/3) ÷ 2
Find:
 (7/3) ÷ 2

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

We have,
= (7/3) × reciprocal of 2
= (7/3) × (1/2)
= (7 × 1) / (3 × 2)
= 7/6
= 1 whole 1/6

Thanks